OpenStudy (anonymous):

Check my work please: dy/dx of Ln[x*y^2]=24 is -2x/y

OpenStudy (anonymous):

dy/dx = -2xy

OpenStudy (anonymous):

i got something really different are you trying to find y' ?

OpenStudy (anonymous):

d/dx (xy^2) = d/dx(y), (x)(2y(dy/dx)) + (y^2)(1) = 0, 2xy(dy/dx + y^2 = 0, 2xy(dy/dx) = -y^2

OpenStudy (anonymous):

oops in the first part, d/dx(xy^2) = d/dx(24)* my bad

OpenStudy (anonymous):

\[\ln (xy^{2}) = 24\]find dy/dx

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

ur answer is -2xy

OpenStudy (anonymous):

what did you do with the natural log Ln ?

OpenStudy (anonymous):

im sry, so stupid

OpenStudy (anonymous):

forgot it

OpenStudy (anonymous):

hehe it happens

OpenStudy (anonymous):

thank you MathMind

OpenStudy (anonymous):

-y/ 2x

OpenStudy (anonymous):

thats the correct answer